A recent broadcast of a television show had a 15 share, meaning that among 6000 monitored households with TV sets in use, 15%p-value also
Answer #1

Solution :

Given that,

n = 6000

\hat p = 0.15

Hypothesis:

H0 : P = 0.25   

Ha : P < 0.25

Test Statistic:

Z =  \hat p - \hat{} \hat{}P / ( \sqrt ((P * (1 - P)) / n)

  Z = 0.15 - 0.25 / ( \sqrt (( 0.25 * 0.75) / 6000)

Z = -17.89

Z = -17.89

P-value = 0.0000

Decision Region : If P-value \leq    \alpha = 0.01 then reject Null hypothesis(H0)

P-value = 0.0000 < \alpha = 0.01, reject H0

There is sufficient evidence to support the claim.

An advertiser that among the households with TV sets in use , less than 25% were tuned into the proggram.